train accelerates from 36 km/h to 72km/h in 20 sec. find the distance travelled by car

Grade:2

2 Answers

Saurabh Koranglekar
3 years ago
Dear student

u = 10 m/s
v= 20 m/s

v^2 = u^2 + 2aS

aS = 150

a=0.5 m/s^2

s=300 m
Jayavardhan
3 years ago
1 kmph = 5/18 m/s
1 m/s = 18/5 m/s
36 kmph= 36×5/18 m/s =10m/s
72 kmph=72×5/18 m/s = 20 m/s
time given = 20 seconds
we know that acceleration=[final velocity-initial velocity]÷time taken = [v-u]÷time taken
so, acceleration = [20m/s – 10m/s]÷20s=0.5m/s2
according to third equation of motion v- u2 = 2as [ where v = final velocity , u = initial velocity , a = acceleration, s = distance]
v2 = 400 m2/s2
u2 = 100 m2/s2
a=0.5
s=?
substitute the valus in third equation of motion,
400m2/s2 – 100m2/s2 = 2 × 5/10 m/s×s
300m = s
 
I HOPE METHOD WILL SOLVE YOUR DOUBT.